# General Exploratory Bonus for Optimistic Exploration in RLHF

### v1

@@ 1 Introduction
To address this, some works (Das et al., 2024; Ji et al., 2024; Mehta et al., 2023; Qi et al., 2025; Liu et al., 2024) have attempted to devise sample-efficient algorithms, inspired by the principle optimism in the face of uncertainty. As illustrated in Fig. 1 (2, top), the principle aims to generate responses with high epistemic uncertainty, thus encouraging data collection in unexplored regions for further training. To operationalize the principle, recent research (Zhang et al., 2024a; Xie et al., 2024; Cen et al., 2025) encourages exploration by adding exploratory bonuses to the reward modeling, which is practically optimizable for large language models. These methods intend to artificially inflate rewards in underexplored regions, nudging the policy toward more informative data collection.
@@
We validate GEB on a large-scale alignment task across different divergences and model backbones. Empirically, GEB consistently yields stronger alignment compared to its counterpart of passive exploration. For example, the three GEB variants that we consider generally outperform the iterative f-DPO (Xiong et al., 2024a) across different divergence regulations, while the most performant variant surpasses several existing optimistic exploration methods that incorporate exploratory bonuses (Zhang et al., 2024a; Xie et al., 2024; Cen et al., 2025). By analyzing the distribution of sampled responses, we validate that GEB can successfully encourage sampling in the region of small $\pi_{\textrm{ref}}$, thereby effectively achieving optimistic exploration.
We summarize our main contributions:
@@
Figure 1: The upper part compares passive exploration and optimistic exploration. Optimistic exploration stimulates the trajectories $\tau$ of small $\pi_{\textrm{ref}}$ (seldom visited/uncertain). While passive exploration sticks to the high-$\pi_{\textrm{ref}}$ region, failing to approach global optima. The lower part contrasts the effect of the exploration bonus term in optimistic reward modeling between prior works and our GEB. Prior works often emphasize rewards in frequently visited regions, which constrains exploration within certain areas. In contrast, our GEB amplifies rewards in seldom-visited regions, thereby encouraging further sampling in uncertain areas and successfully achieving optimistic exploration.
@@ 2 Preliminaries
Iterative online RLHF. The effectiveness of iterative online RLHF (Xiong et al., 2024b; Dong et al., 2024) has been validated in many real-world systems such as Claude (Bai et al., 2022) and LLaMA-series (Touvron et al., 2023; Grattafiori et al., 2024). The algorithm proceeds in $T$ rounds, with each round having two steps: (1) the $\pi_{t}$ is learned with the current dataset $\mathcal{D}_{t}$, and then samples $x\sim\rho,\ (y_{1},y_{2})\sim\pi_{t}(\cdot|x)$; (2) Human evaluator annotate the preference of $(x,y_{1},y_{2})\rightarrow(x,y^{w},y^{l})$ to form $\mathcal{D}_{t+1}$, where the prompt $x$ is sampled from an independent distribution $\rho$, response $y_{1},y_{2}$ are two response sampled from the policy of the $t$-th iteration $\pi_{t}$. When computing $\pi_{t}$ with dataset $D_{t}$ in the step (1) of each iteration, a reward function $r_{t}(x,y)$ is first learned from a collected human preference data $\mathcal{D}_{t}=\{(x,y^{w},y^{l})\}$, where $y^{w},y^{l}$ denote the preferred and dispreferred response to $x$, respectively. Reward modeling typically follows the Bradley-Terry objective (Bradley & Terry, 1952):
@@
where $\sigma$ denotes the sigmoid function. Given the learned reward function $r_{t}$, the policy $\pi_{t}$ is then updated to maximize the expected reward, often with a KL-regularization as follows
@@
where $\beta>0$ is a hyperparameter, and $\pi_{\textrm{ref}}$ is the reference model.
Sample inefficiency of iterative online RLHF. Online sampling for standard online RLHF algorithms is carried out passively, relying solely on the inherent randomness of the LLM policy. However, if the policy places a small probability mass on the optimal action, passive exploration may fail to ever explore this action. Theoretical analyses (Muldrew et al., 2024; Dong et al., 2024) and empirical evidence (Chen et al., 2025; Cen et al., 2025) present that the passive approach fails to sufficiently explore the prompt-response space. Particularly, Xie et al. (2024) demonstrate that the sample complexity can be exponential in $1/\beta$ for passive exploration, which is unacceptable in the small-$\beta$ regime. Therefore, several works (Xie et al., 2024; Cen et al., 2025; Zhang et al., 2024a) propose exploratory bonuses to implement optimistic exploration for efficient sampling. However, in the next section, we will show that prior formulations cannot provably achieve optimism.
@@ 3 Exploratory Bonus and How It Can Fail
In this section, we will first provide the iterative online RLHF formulation with an exploratory bonus (Section 3.1). We then theoretically prove that the existing formulation can fail to achieve optimistic exploration under both KL-constrained RLHF (Section 3.2) and a more general $\alpha$-divergence-regularized RLHF (Section 3.3), motivating our proposed method in Section 4.
@@ 3.1 Exploratory Bonus
To improve the sample efficiency of iterative online RLHF, recent works (Zhang et al., 2024a; Cen et al., 2025) introduce exploratory bonuses, which try to encourage optimistic exploration. These approaches modify the standard loop by adding an exploratory bonus term $\mathcal{L}_{\text{bonus}}$. Specifically, in the $t$-th iteration, the reward model $r_{t}$ and policy $\pi_{t}$ are optimized by
@@
To achieve optimistic exploration, the bonus term is expected to stimulate probability increase more prominently in unexplored areas. Formally, we have the following definition:
@@
When a reward model $r$ and a policy $\pi$ is computed with Eq. 3 and Eq. 4, the exploratory bonus $\mathcal{L}_{\text{bonus}}(r)$ achieves optimism, if
$\frac{\partial^{2}\mathcal{L}_{\text{bonus}}(r)}{\partial\pi(y|x)\partial\pi_{s}(y|x)}<0,$ (5)
where $\pi_{s}(y|x)$ is an ideal sampling distribution for response at the current iteration.
To interpret the optimism condition of the exploration bonus term in Definition. 3.1, we consider the policy-reparameterized reward model $r(x,y):=r(\pi)$, which can be derived from the closed-form solution of Eq. 2 as $\pi(y|x)=\frac{\exp(r(x,y)/\beta)}{Z(x)}$, where $Z(x)=\mathbb{E}_{y\sim\pi_{\textrm{ref}}}\exp(r(x,y)/\beta)$ is a normalization function. This yields the reward model expressed via the policy (Rafailov et al., 2024):
$r(\pi)=\beta\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}+\log Z(x).$
@@
Eq. 5 requires the gradient $\partial\mathcal{L}_{\text{bonus}}(r(\pi))/\partial\pi(y|x)$ to be negatively correlated with the sampling probability $\pi_{s}$. In other words, responses rarely sampled under $\pi_{s}$ (i.e., uncertain or underexplored outputs) should receive a larger ascending of the policy distribution $\pi$, i.e., larger $\partial\mathcal{L}_{\text{bonus}}(r(\pi))/\partial\pi(y|x)$. In practice, $\pi_{s}$ can be substituted by the reference model or intermediate checkpoints. We adopt the commonly used $\pi_{\textrm{ref}}$ as $\pi_{s}$ in our following demonstration.
@@ 3.2 Failure Under KL-constrained RLHF
Previous works, including Zhang et al. (2024a) and Cen et al. (2025), formulate the exploratory bonus with $\mathcal{L}_{\text{bonus}}(r)=\max_{\pi}\mathcal{J}_{\beta,KL}(\pi,r)$. In this case, optimizing exploratory bonus in Eq. 3 becomes a min-max bi-level objective as $\min_{r}-\kappa\max_{\pi}[\mathbb{E}_{x,y\sim\pi}r(x,y)-\beta\mathbb{D}_{\text{KL}}(\pi\|\pi_{\textrm{ref}})]$. Intuitively, they intend to make $r$ not only fit the observed data by $\mathcal{L}_{BT}$, but also have a larger reward in unobserved regions by maximizing the $\max_{\pi}\mathbb{E}_{x,y\sim\pi}r(x,y)$ in $\mathcal{L}_{\text{bonus}}(r)$. Here, we theoretically show that such formulations can suffer from optimism failures under KL-regularized RLHF.
@@
Proof  First, the inner maximization of the bonus term admits a closed-form solution, $\pi^{*}(y|x)=\pi_{\textrm{ref}}(y|x)e^{\frac{r(x,y)}{\beta}}/Z(x)$ where $Z(x)=\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}e^{\frac{r(x,y)}{\beta}}$ is a normalization term. Substituting this solution reduces the bi-level training objective of $r_{2}$ to a single-level form:
As shown in Rafailov et al. (2024), the log-ratio $\beta\log\pi_{\theta}(y|x)-\beta\log\pi_{\textrm{ref}}(y|x)$ represents the same class of the original reward function $r$ through Eq. 4, thus the reward $r_{1},r_{2}$ can be reparameterized by the log-ratio. Plugging this into Eq. 6 yields
Since the second term equals 0, the reparameterized Eq. 6 is exactly the vanilla DPO loss, which is the same as the reparameterized training objective of $r_{1}$. Thus, the exploratory bonus in the reward training objective has no effect on the final policy set. $\square$
The lemma proves that incorporating the exploratory bonus $\mathcal{L}_{\text{bonus}}(r)=\max_{\pi}\mathcal{J}_{\beta,\text{KL}}(\pi,r)$ into the reward training objective fails to induce the policy model to sample from low-$\pi_{\textrm{ref}}(y|x)$ regions, i.e., unexplored responses. In other words, the bonus term is ineffective at inducing optimism. We next extend the result beyond KL regularization to a more general class of $\alpha$-divergence family.
@@ 3.3 Generalization to $\alpha$-divergence-constrained RLHF
In this subsection, we theoretically show that the failure of optimism can broadly extend to the $\alpha$-divergence class. Many common divergences, such as reverse KL-divergence, Hellinger distance, and forward KL-divergence, are special cases of $\alpha$-divergence.
@@
Consider objective $\mathcal{J}_{\beta,f}(\pi,r)=\mathbb{E}_{x\sim\rho,y\sim\pi(y|x)}r(x,y)+\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(y|x)}f(\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})$, where $f$ belongs to $\alpha$-divergence class. If a reward is trained with $r=\arg\min_{r}[\mathcal{L}_{BT}(\mathcal{D},r)-\kappa\mathcal{L}_{\text{bonus}}]$ and a policy $\pi$ is updated by $\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ with $\mathcal{L}_{\text{bonus}}=\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$, the gradient of the bonus satisfies $\frac{\partial^{2}\mathcal{L}_{\text{bonus}}(r(\pi))}{\partial\pi\partial\pi_{\textrm{ref}}}\geq 0$, which means $\mathcal{L}_{\text{bonus}}$ encourage trajectories with large $\pi_{\textrm{ref}}$ more strongly, in contradiction to the optimism principle (Definition 3.1).
@@
$\pi_{f}^{*}=\frac{1}{Z(x)}\pi_{\textrm{ref}}(y|x)(f^{\prime})^{-1}(r/\beta),\quad r(x,y)=\beta f^{\prime}(\frac{\pi^{*}(y|x)}{\pi_{\textrm{ref}}(y|x)}Z(x)),$ (8)
where $Z(x)$ is a normalization term and $(f^{\prime})^{-1}$ is the inverse function of $f^{\prime}$. The bi-level objective can be similarly transformed to a single level one by canceling the inner maximization $\max_{\pi}$ by Eq. 8. The single-level objective can be written as $r_{t}=\arg\min_{r}\mathcal{L}_{BT}(\mathcal{D},r)-\kappa\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{1}{Z(x)}(f^{\prime})^{-1}(\frac{r(x,y)}{\beta})\cdot r(x,y)-\beta f(\frac{1}{Z(x)}(f^{\prime})^{-1}(\frac{r(x,y)}{\beta})$. Since the policy is computed by $\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$, the reward can be reparameterized by the policy with Eq. 8, which fortunately cancels $Z(x)$. Then the optimistic reward-modeling objective can be reparameterized as
@@
$f$ exploratory bonus
reverse KL constant
forward KL $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(y|x)}\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}$
Hellinger distance $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(y|x)}\sqrt{\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}}$
According to Lemma. 3.2, we enumerate several exploratory bonus under different $\alpha$-divergence in Table 1. The listed bonuses are simplified by removing constant coefficients and bias. The listed exploratory bonuses generally force the policy model to maximize the possibility of trajectories sampled by the reference model, not the underexplored ones. We further prove that it actually drives $\pi$ to collapse toward $\pi_{\textrm{ref}}$ and that the failure extends beyond $\alpha$-divergence to other $f$-divergences.
@@
When $f$ belongs to $f$-divergence, and the reward function is obtained by $\hat{r}=\arg\min_{r}[\mathcal{L}_{BT}(\mathcal{D}_{t},r)-\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)]$ and the policy is updated by $\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r_{t})$, the bonus term $\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ induces the policy model $\pi$ to coincide with $\pi_{\textrm{ref}}$ when $xf^{\prime\prime}(x)$ is a monotone function.
The detailed proofs are in Appendix B.1. The monotone increase of $xf^{\prime\prime}(x)$ can be satisfied by a broader divergence class besides $\alpha$-divergence, including JS-divergence and Pearson $\chi^{2}$.
@@
Optimization of exploratory bonus in Eq. 3 is a min-max bi-level objective as $\min_{r}-\kappa\max_{\pi}[\mathbb{E}_{x,y\sim\pi}r(x,y)-\beta\mathbb{D}_{\text{KL}}(\pi\|\pi_{\textrm{ref}})]$. Due to the inner maximization $\max_{\pi}$, the divergence constraint implicitly makes $\pi$ close to $\pi_{\textrm{ref}}$ to avoid the divergence penalty. Considering the outer minimization $\min_{r}$, $r$ is forced to provide large rewards on region of high $\pi$ to maximize the reward expectation. Their combination implicitly makes $r$ focus more on region of large $\pi_{\textrm{ref}}$. Since samples with large $\pi_{\textrm{ref}}$ is easily rolled out from scratch, previous exploratory bonuses merely concentrate sampling on regions that are already easy to explore, contradictory to the optimism principle, which requires encouraging responses $y$ rarely sampled by the reference model.

### v4

@@ 1 Introduction
To address this, some works (Das et al., 2024; Ji et al., 2024; Mehta et al., 2023; Qi et al., 2025; Liu et al., 2024) have attempted to devise sample-efficient algorithms, inspired by the principle optimism in the face of uncertainty. As illustrated in Fig. 1 (2, top), the principle aims to generate responses for regions of high epistemic uncertainty, thus encouraging data collection in unexplored areas for further training. To operationalize this, recent attempts (Zhang et al., 2024a; Xie et al., 2024; Cen et al., 2025) encourage exploration by adding exploratory bonuses to reward modeling, which is practically optimizeable for large language models. These methods intend to artificially inflate rewards in underexplored regions, nudging the policy toward more informative data collection.
@@
We validate GEB on a large-scale alignment task across different divergences and model backbones. Empirically, GEB consistently yields stronger alignment compared to its counterpart of passive exploration. For example, the three GEB variants that we consider generally outperform the iterative f-DPO (Xiong et al., 2024a) across different divergence regulations, while the most performant variant surpasses several existing optimistic exploration methods that incorporate exploratory bonuses (Zhang et al., 2024a; Xie et al., 2024; Cen et al., 2025). By analyzing the distribution of sampled responses, we validate that GEB can successfully encourage sampling in the region of small $\pi_{\textrm{ref}}$, thereby effectively achieving optimistic exploration. We summarize our main contributions:
Figure 1: The upper part compares passive exploration and optimistic exploration. Optimistic exploration stimulates the trajectories $\tau$ of small $\pi_{\text{ref}}$ (seldom visited/uncertain), while passive exploration sticks to the high-$\pi_{\text{ref}}$ region, failing to approach global optima. The dashed line separates regions of high vs. low likelihood under the learning policy $\pi_{\theta}$. The lower part contrasts the effect of the exploration bonus term in optimistic reward modeling between prior works and our GEB. Prior works often emphasize rewards in frequently visited regions, which constrains exploration within certain areas. In contrast, our GEB amplifies rewards in seldom-visited regions, thereby encouraging further sampling in uncertain areas and successfully achieving optimistic exploration.
@@ 2 Preliminaries
Iterative online RLHF.
[…]
@@
where $\sigma$ denotes the sigmoid function. Next, in each step (ii), given the learned reward function $r_{t}$, the policy $\pi_{t}$ is updated to maximize the expected reward, often with a KL-regularization as follows
@@
where $\beta>0$ is a hyperparameter and $\pi_{\textrm{ref}}$ is the reference model. The effectiveness of iterative online RLHF (Xiong et al., 2024b; Dong et al., 2024) has been validated in various real-world systems such as Claude (Bai et al., 2022) and LLaMA-series (Touvron et al., 2023; Grattafiori et al., 2024), but there is still much room for improvement in terms of sample-efficient exploration.
Sample inefficiency of iterative online RLHF.
In online RLHF, standard online sampling is usually performed passively, relying solely on the LLM policy’s inherent randomness. However, if the policy assigns a small probability to the optimal action, passive exploration may never explore it. Some recent theoretical analyses (Muldrew et al., 2024; Dong et al., 2024) and empirical evidence (Chen et al., 2025; Cen et al., 2025) present that the passive approach fails to sufficiently explore the prompt-response space. Particularly, Xie et al. (2024) demonstrates that the sample complexity can be exponential in $1/\beta$ for passive exploration, which is unacceptable in the small-$\beta$ regime. After then, follow-up studies propose to implement the principle “optimism towards uncertainty” into RLHF algorithms, i.e., encourage exploration of uncertain trajectories. Several works on this try to estimate uncertainty by leveraging some uncertainty quantification techniques, such as elliptical potential (Bai et al., 2022), Bayesian modeling (Qi et al., 2025), and epistemic neural network training (Liu et al., 2024). However, these methods are generally computationally prohibitive in LLM-scale settings. Therefore, recent works (Xie et al., 2024; Cen et al., 2025; Zhang et al., 2024a) propose exploratory bonuses for optimistic exploration, which can be computationally more tractable for LLM-based optimization.
@@ 3 Exploratory Bonus and How It Can Fail
In this section, we first provide the iterative online RLHF formulation with an exploratory bonus (Section 3.1). We then theoretically prove that the existing formulation can fail to achieve optimistic exploration under both KL-constrained RLHF (Section 3.2) and a more general $\alpha$-divergence-regularized RLHF (Section 3.3), motivating our proposed method in Section 4.
@@ 3.1 Exploratory Bonus
To improve the sample efficiency of iterative online RLHF, recent works (Zhang et al., 2024a; Cen et al., 2025) introduce exploratory bonuses, which aim to encourage the policy model to explore the under-visited space given an optimistic reward estimation. These approaches modify the standard RLHF loop by adding an exploratory bonus term $\mathcal{L}_{\text{bonus}}$ in the reward modeling phase. Specifically, in the $t$-th iteration, the reward model $r_{t}$ and policy $\pi_{t}$ are optimized by
@@
To boost exploration, the bonus term is designed to amplify the probability mass of policy more in underexplored areas rather than incentivizing it solely towards high empirical reward areas. As mentioned in § 2, early works on RLHF optimistic exploration are computationally prohibitive in the LLM fine-tuning regime. Thus, it is necessary to set a new approach that not only aligns with the principle of optimism in the face of uncertainty but is also cost-effective. For this, we derive a new condition for the exploration bonus to achieve optimism, avoiding direct uncertainty quantification:
@@
Given an input prompt $x$ and a response $y$, when a reward model $r$ and a policy $\pi$ are computed with Eq. 3 and Eq. 4, respectively, the exploratory bonus $\mathcal{L}_{\text{bonus}}$ achieves optimism, if
$\frac{\partial}{\partial\pi_{s}(y|x)}(\frac{\partial\mathcal{L}_{\text{bonus}}(r(x,y))}{\partial\pi(y|x)})<0,$ (5)
where $\pi_{s}$ is a typical sampling policy, a joint policy on all iterations up to the current iteration.
Specifically, at the $t$-th iteration, the typical sampling policy $\pi_{s}=\pi_{1}\circ\pi_{2}\circ\cdots\circ\pi_{t}$ is a joint distribution of all previous policies up to the current iteration. This distribution is not directly computable; rather, it serves as a theoretical construct describing how responses in $\mathcal{D}_{t}$ are generated. In Eq. 5, rather than characterizing $\mathcal{L}_{\text{bonus}}$ in its original function space, we define it by a condition on its partial derivatives with respect to two policies: current policy and typical policy. This new optimism condition not only enables us to flexibly define $\mathcal{L}_{\text{bonus}}$, but also serves as a core tool for theoretically analyzing existing methods. Although $\mathcal{L}_{\text{bonus}}$ appears unrelated to the current policy $\pi$—as it is defined in terms of the reward model r(x,y)r(x,y)the policy-reparameterized reward $r_{\pi}(x,y)$ allows us to express $r(x,y)$ directly in terms of $\pi$ as follows (Rafailov et al., 2024):
$r(x,y):=r_{\pi}(x,y)=\beta\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}+\beta\log Z(x),~~~~\text{where}~~Z(x)=\mathbb{E}_{y\sim\pi_{\textrm{ref}}}\exp(r(x,y)/\beta).$
This is derived from the closed-form solution of the proximal preference optimization (Eq. 2) as $\pi(y|x)=\frac{\exp(r(x,y)/\beta)}{Z(x)}$. Thanks to this alternative representation of the reward, we can express and interpret the bonus term with respect to our current policy $\pi$, which yields the following implication.
@@
Our new optimism condition requires the derivative of the bonus term with respect to the current policy, $\partial\mathcal{L}_{\text{bonus}}(r_{\pi}(x,y))/\partial\pi(y|x)$, to be negatively correlated with the typical policy $\pi_{s}$. In other words, as a response $y$ is likely rare sample under $\pi_{s}$ (i.e., uncertain or underexplored responses), it should receive a larger ascending force in the policy distribution $\pi$, i.e., a higher $\partial\mathcal{L}_{\text{bonus}}(r_{\pi}(x,y))/\partial\pi(y|x)$. This new definition of the optimism principle, which is specified through the lens of partial derivative alignment, ensures the exploratory bonus nudges the exploration towards an uncertain response region without explicit uncertainty quantification. In practice, $\pi_{s}$ can be substituted by the reference model policy $\pi_{\textrm{ref}}$ or intermediate checkpoints of $\pi$ across iterations. We adopt the commonly used $\pi_{\textrm{ref}}$ as $\pi_{s}$ in our following demonstration.
@@ 3.2 Failure Under KL-constrained RLHF
Previous works, including Zhang et al. (2024a) and Cen et al. (2025), formulate the exploratory bonus with $\mathcal{L}_{\text{bonus}}(r)=\max_{\pi}\mathcal{J}_{\beta,KL}(\pi,r)$. Under this formulation, optimizing the exploratory bonus in Eq. 3 yields a min–max bilevel objective: $\min_{r}-\kappa\max_{\pi}[\mathbb{E}_{x\sim\rho,y\sim\pi}r(x,y)-\beta\mathbb{D}_{\text{KL}}(\pi\|\pi_{\textrm{ref}})]$. Intuitively, this objective encourages $r$ not only to fit the observed data via $\mathcal{L}_{BT}$ but also to assign high reward to unobserved regions by maximizing $\max_{\pi}\mathbb{E}_{x\sim\rho,y\sim\pi}r(x,y)$ in $\mathcal{L}_{\text{bonus}}(r)$. Here, we theoretically show that such formulations can suffer from optimism failures under KL-regularized RLHF.
@@
See the proof in Appendix B.1. The lemma shows that incorporating the exploratory bonus $\mathcal{L}_{\text{bonus}}(r)=\max_{\pi}\mathcal{J}_{\beta,\text{KL}}(\pi,r)$ into the reward training objective fails to induce the policy to sample from low-$\pi_{\textrm{ref}}(y|x)$ regions, i.e., unexplored responses. That is, $\mathcal{L}_{\text{bonus}}$ is ineffective for optimism. We next extend the result beyond KL divergence to a more general class of $\alpha$-divergence families.
@@ 3.3 Generalization to $\alpha$-divergence-constrained RLHF
$f$ exploratory bonus
reverse KL constant
forward KL $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}$
Hellinger distance $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\sqrt{\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}}$
In this subsection, we theoretically show that the failure of optimism can broadly be extended to the $\alpha$-divergence class. Many common divergences, such as reverse KL-divergence, Hellinger distance, and forward KL-divergence, are special cases of $\alpha$-divergence.
@@
[…]
@@
$\pi_{f}^{*}(y|x)=\frac{1}{Z(x)}\pi_{\textrm{ref}}(y|x)(f^{\prime})^{-1}(r(x,y)/\beta),\quad r_{\pi}(x,y)=\beta f^{\prime}(\frac{\pi^{*}(y|x)}{\pi_{\textrm{ref}}(y|x)}Z(x)),$ (6)
where $Z(x)$ is a normalization term and $(f^{\prime})^{-1}$ is the inverse function of $f^{\prime}$. The bi-level objective can be similarly transformed to a single level one by canceling the inner maximization $\max_{\pi}$ by Eq. 6. The single-level objective can be written as $r_{t}=\arg\min_{r}\mathcal{L}_{BT}(\mathcal{D},r)-\kappa\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{1}{Z(x)}(f^{\prime})^{-1}(\frac{r(x,y)}{\beta})\cdot r(x,y)-\beta f(\frac{1}{Z(x)}(f^{\prime})^{-1}(\frac{r(x,y)}{\beta})$. Since the policy is computed by $\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$, the reward can be reparameterized by the policy with Eq. 6, which fortunately cancels $Z(x)$. Then, the optimistic reward-modeling objective can be reparameterized as
@@
According to Lemma 3.2, we summarize several forms of exploratory bonus induced by different $\alpha$-divergences in Table 1. For clarity, these expressions are presented after removing constant coefficients and additive biases. In every case, the resulting bonus encourages the policy to place more probability mass on responses that the reference model already samples frequently, rather than on underexplored responses. Now, we further prove that it actually drives $\pi$ to collapse toward $\pi_{\textrm{ref}}$ and that the failure extends beyond $\alpha$-divergence to other $f$-divergences.
@@
[…]
The monotonic increase of $xf^{\prime\prime}(x)$ can be satisfied by a broader divergence class beyond $\alpha$-divergence, including JS-divergence and Pearson $\chi^{2}$. Please see the detailed proofs in Appendix B.2.
@@
The optimization of the exploratory bonus in Eq. 3 is a min-max bi-level objective, $\min_{r}-\kappa\max_{\pi}[\mathbb{E}_{x\sim\rho,y\sim\pi}r(x,y)-\beta\mathbb{D}_{\text{KL}}(\pi\|\pi_{\textrm{ref}})]$. Due to inner maximization $\max_{\pi}$, the divergence constraint implicitly makes $\pi$ close to $\pi_{\textrm{ref}}$ to reduce the KL divergence. Meanwhile, outer minimization $\min_{r}$ forces $r$ to provide high rewards in the region of high $\pi$ to maximize the expected reward. Their combination implicitly makes $r$ focus more on the region of high $\pi_{\textrm{ref}}$. As responses in high $\pi_{\textrm{ref}}$ area are easily sampled from scratch, prior exploratory bonuses just concentrate sampling on regions that are already frequently visited, contradictory to the optimism principle, which requires encouraging exploration for responses $y$ rarely sampled by the reference model.